OpenStudy (anonymous):

Two Particles are projected vertically upwards with speed 20m/s and 100m/s from a top of a tower if their time so flight are in ratio 1:3 find the height of the tower?

OpenStudy (anonymous):

use second equation of motion separately for both..to find t1 and t2 in terms of h...using g=-10,s=-h,u1=20,u2=100. use the ratio condition to find the value of h..

OpenStudy (anonymous):

calculated it??

OpenStudy (anonymous):

k.can u show me where u got struck? anyone else calculated it?

OpenStudy (anonymous):

No not yet....

OpenStudy (anonymous):

Let height of tower be h. Therefore time for the first projectile = u = +20 height = -h.Let time be t1. -h = +20t1 - 5t1^2 (As g = -10)-----Eq(1) For second particle. u = +100 height = -h. Let time be t2. - h = 100t2 - 5t2^2. ---------Eq(2) Eq(1) = Eq(2) ( -h = -h) 100t2 - 5t2^2 = 20t1 - 5t1^2. Now t2 = 3t1. 300t1 - 5(9t1^2) = 20t1 - 5t1^2 40t1^2 - 280t1 = 0 Since t1 is not 0. => 40t1 = 280 t1 = 7 seconds. t2 = 21 seconds Satisfying t1 = 7 seconds in eq(1) -h = 20*7 - 5*49 => h = 105m

OpenStudy (anonymous):

Thxxx but how h1=h2

OpenStudy (anonymous):

it is the final displacement of both balls..

OpenStudy (anonymous):

@open_study1 the tower is the same for both the particles so h1 will be equal to h2